Consider this molecular-level representation of a gas. co O Play If the partial pressure of the diatomic gas is 0.390 atm, what is the total pressure? Number atm
Answer #1

Balanced equation:
2 Al(s) + 6 HCl ====> 2 AlCl3(aq) + 3 H2(g)

Reaction type: single replacement

2.9 gm of Aluminum = 2.9 /26.98 = 0.10748 Moles

0.10748 Moles of Aluminum will produce 0.1612 Moles (0.10748 x 3/2) of hydrogen

We can find the volume of Hydrgen gas at STP using ideal gas equation

PV = nRT

where P is the absolute pressure, V is the volume of gas , n is the amount of gas (SI unit moles),, and T is the Temperature .

1 x V = 0.1612 x 0.0821 x 273

V = 3.61 L

Volume of Hydrogen gas 3.61 Liter


answered by: ANURANJAN SARSAM
Answer #2

Balanced equation:
2 Al(s) + 6 HCl ====> 2 AlCl3(aq) + 3 H2(g)

Reaction type: single replacement

2.9 gm of Aluminum = 2.9 /26.98 = 0.10748 Moles

0.10748 Moles of Aluminum will produce 0.1612 Moles (0.10748 x 3/2) of hydrogen

We can find the volume of Hydrgen gas at STP using ideal gas equation

PV = nRT

where P is the absolute pressure, V is the volume of gas , n is the amount of gas (SI unit moles),, and T is the Temperature .

1 x V = 0.1612 x 0.0821 x 273

V = 3.61 L

Volume of Hydrogen gas 3.61 Liter